7. Suppose X and Y have joint pdf f(x,y) = 24x y if x >0, y>0,x+y<1 and 0 otherwise. Find P(Y > 2x).
Answer #1

922x = 9-2x70 - uty=1 solving Y=20 and x+y=1 X444 1 x+ 2X=1 24...496) 30=1 x=13 ->y=2017 .: point of intersection in G 1 11-Thumbsup 一 324 txel |- 华亭一年。 -- | c1= - -- - 1